# Faster Convergence of Riemannian Stochastic Gradient Descent with Increasing Batch Size

### v3

@@ Abstract
We have theoretically analyzed the use of Riemannian stochastic gradient descent (RSGD) and found that using an increasing batch size leads to faster RSGD convergence rate than using a constant batch size not only with a constant learning rate but also with a decaying learning rate, such as cosine annealing decay and polynomial decay. The convergence rate of RSGD improves from $O(\sqrt{T^{-1}+\text{const.}})$ with a constant batch size to $O(T^{-\frac{1}{2}})$ with an increasing batch size, where $T$ denotes the number of iterations. Using principal component analysis and low-rank matrix completion tasks, we investigated, both theoretically and numerically, how increasing batch size affects computational time as measured by stochastic first-order oracle (SFO) complexity. Increasing batch size reduces the SFO complexity of RSGD. Furthermore, our numerical results demonstrated that increasing batch size offers the advantages of both small and large constant batch sizes.
@@ 1 Introduction
Stochastic gradient descent (SGD) Robbins & Monro (1951) is a basic algorithm, which is widely used in machine learning Liu et al. (2021); He et al. (2016); Krizhevsky et al. (2012). Riemannian stochastic gradient descent (RSGD) was proposed in Bonnabel (2013). Riemannian optimization Absil et al. (2008); Boumal (2023), which studies RSGD and its derivations, has been attracting attention for its use including many machine learning tasks. For example, it has been used in principal components analysis Breloy et al. (2021); Liu & Boumal (2020); Roy et al. (2018), low-rank matrix completion Vandereycken (2013); Nguyen et al. (2019); Boumal & Absil (2015); Kasai & Mishra (2016), convolutional neural networks Wang et al. (2020); Huang et al. (2017); Chen et al. (2017), graph neural networks Zhu et al. (2020); Chami et al. (2019); Liu et al. (2019), and has been used in applications of optimal transportation theory Lin et al. (2020b); Weber & Sra (2023).
It is well established that the performance of Euclidean SGD strongly depends on both the batch size (BS) and learning rate (LR) settings Goyal et al. (2017); Smith et al. (2018); Zhang et al. (2019); Lin et al. (2020a). In the context of Riemannian SGD, Ji et al. (2023); Bonnabel (2013); Kasai et al. (2019; 2018) addressed the use of a constant batch size and Sakai & Iiduka (2025); Han & Gao (2022) addressed the use of an adaptive BS (although the latter did not consider RSGD specifically). Additionally, Umeda & Iiduka (2024) found that increasing BS accelerated the convergence of Euclidean SGD. In parallel, various decaying LR strategies, including cosine annealing Loshchilov & Hutter (2017) and polynomial decay Chen et al. (2018), have been devised and their effectiveness has been demonstrated in the Euclidean case. Motivated by these findings, we have developed an approach that achieves a faster convergence rate for RSGD by using an increasing BS in combination with several types of a decaying LR schedules, including ones that have not yet been analyzed in the Riemannian context. Our numerical results demonstrate that an increasing BS leads to faster convergence in terms of both steps and computational time.
@@ 1.1 Previous Results and Our Contributions
Reference and Theorem Batch Size Learning Rate Convergence Analysis
Bonnabel (2013) s.a.r. $\underset{t\to\infty}{\lim}\mathbb{E}[\mathrm{grad}f(x_{t})]=0\text{ a.s.}$
Zhang & Sra (2016) $\eta_{t}=O(\frac{1}{L_{r}+\sqrt{t}})$ $\mathbb{E}[f(\bar{x}_{T})-f^{\star}]=O(\frac{\sqrt{T}}{C+T})$
Tripuraneni et al. (2018) Constant $\eta_{t}=O(\frac{1}{t^{\alpha}})$ $\mathbb{E}\|\tilde{\Delta}_{T}\|=O(\frac{1}{\sqrt{T}})$
Hosseini & Sra (2020) Constant $\eta_{t}=O(\frac{1}{\sqrt{T}})$ $\|G_{T}\|=O(\frac{1}{\sqrt[4]{T}})$
Sakai & Iiduka (2024) Constant $\eta_{t}=\frac{1}{\sqrt{t+1}}$ $\|G_{T}\|=O(\frac{\sqrt{\log T}}{\sqrt[4]{T}})$
Sakai & Iiduka (2025) Increasing Constant $\|G_{T}\|=O(\frac{1}{\sqrt{T}})$
Theorem 1 Constant Constant and Decay $\|G_{T}\|=O(\sqrt{\frac{1}{T}+C})$
Theorem 2 Increasing Constant and Decay $\|G_{T}\|=O(\frac{1}{\sqrt{T}})$
Theorem 3 Increasing Warm-up $\|G_{T}\|=O(\frac{1}{\sqrt{T-T_{w}}})$
Theorem 4 Constant Warm-up $\|G_{T}\|=O(\sqrt{\frac{1}{T-T_{w}}+C})$
Table 1: Comparison of RSGD convergence analyses, where $\alpha\in(\frac{1}{2},1),C_{1}\in(0,1),C_{2}$ and $C$ are constants. For $\tilde{x}_{t}\coloneq R_{\tilde{x}_{t-1}}(\frac{1}{t}R_{\tilde{x}_{t-1}}^{-1}(x_{t}))$, $\tilde{\Delta}_{t}\coloneq R_{x^{\star}}^{-1}(\tilde{x}_{t})$. The usual stochastic approximation learning rate rule (s.a.r.) is defined as $\sum_{t=0}^{\infty}\eta_{t}=\infty,\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty$. $T$ is the total number of iterations, and $T_{w}$ is the number of warm-up iterations (see Section 3.3). $\min_{t\in\{0,\cdots,T-1\}}\mathbb{E}[\|\mathrm{grad}f(x_{t})\|_{x_{t}}]$ is represented by $\|G_{T}\|$. ‘Decay’ includes diminishing LR.
@@
Constraint: Bonnabel (2013); Zhang & Sra (2016); Durmus et al. (2021); Sakai & Iiduka (2024) considered Hadamard manifolds or their submanifolds. Tripuraneni et al. (2018) treated Riemannian manifolds that satisfy the condition required to ensure a unique geodesic connecting any two points, including Hadamard manifolds. Sakai & Iiduka (2025) addressed embedded submanifolds of Euclidean space. In contrast, we considered general Riemannian manifolds that encompass all of the above.
@@
Learning rate: Bonnabel (2013) considered LRs satisfying $\sum_{t=0}^{\infty}\eta_{t}=\infty,\sum_{t=0}^{\infty}\eta^{2}_{t}<\infty$, and Zhang & Sra (2016); Tripuraneni et al. (2018); Sakai & Iiduka (2024) considered diminishing LRs of the form $\eta_{t}=\frac{1}{t^{a}}$. Hosseini & Sra (2020); Durmus et al. (2021) considered constant LRs. Sakai & Iiduka (2025) used both constant and diminishing LRs. We considered constant, diminishing, cosine annealing, and polynomial decay LRs, as well as these with warm-up.
@@
Batch size: Since Bonnabel (2013) considered expected risk minimization using expectation instead of sample mean—as is customary in empirical risk minimization—the concept of batch size is not generally applicable in that context. Zhang & Sra (2016); Tripuraneni et al. (2018); Hosseini & Sra (2020); Durmus et al. (2021); Sakai & Iiduka (2024) considered a constant BS. Sakai & Iiduka (2025) and the present work considered an increasing BS. Furthermore, we numerically and theoretically compared using a constant BS and an increasing BS by using stochastic first-order oracle (SFO) complexity. Our contributions are as follows: (i) we theoretically showed that using an increasing BS yields a more favorable SFO rate of $O(\epsilon^{-2})$, whereas using a constant BS with the critical batch size yields a rate of $O(\epsilon^{-4})$; (ii) we numerically observed that increasing BS achieved both a better minimizer—in terms of attaining a smaller gradient norm—and a shorter computational time than using either a small or large constant BS (see Section 4.3 for more details).
@@
Objective function: Bonnabel (2013) considerd three times continuously differentiable, with both the gradient and Hessian uniformly bounded. Tripuraneni et al. (2018) treated functions that are twice continuously differential, subject to additional conditions on the Hessian. Other studies considered only once continuously differentiability. Zhang & Sra (2016) addressed convex and smooth functions, while Durmus et al. (2021) considered strongly convex and smooth functions. Hosseini & Sra (2020); Sakai & Iiduka (2024; 2025) treated more general classes of functions, specifically nonconvex functions with bounded gradients. We treated nonconvex functions without bounded gradients.
@@
Convergence rate: A comparison of the previous and current RSGD convergence analyses is presented in Table 1; here we outline the key results. Tripuraneni et al. (2018) obtained a convergence rate of $O(\frac{1}{\sqrt{T}})$ (with a criterion other than the gradient norm) using a learning rate of $\eta=O(\frac{1}{\sqrt{T}})$; however, this rate decays to zero as $T\to\infty$, making it unsuitable for practical use. Durmus et al. (2021) attained exponential convergence under strong convexity and smoothness assumptions. Sakai & Iiduka (2025) attained a rate of $O(\frac{1}{\sqrt{T}})$ for several adaptive methods, including RSGD. In contrast, our work differs from Sakai & Iiduka (2025) in four key aspects : (i) we provide both theoretical and numerical evidence—based on SFO complexity—that using an increasing batch size leads to better performance than using a fixed batch size; (ii) in addition to a constant LR and diminishing LR, we addressed cosine annealing, polynomial decay, and warm-up LRs; (iii) Sakai & Iiduka (2024) focused on embedded submanifolds of Euclidean space unlike us; (iv) we do not assume bounded gradient norms, thus expanding the range of applicable scenarios.
Summary:
@@
  • We developed a convergence analysis of RSGD incorporating increasing BS, cosine annealing LR, and polynomial decay LR, under assumptions that are more general than those used in prior work. Our analysis demonstrates that using an increasing BS improves the convergence rate of RSGD from $O(\sqrt{\frac{1}{T}+C})$ to $O(\frac{1}{\sqrt{T}})$.
@@
  • We numerically and theoretically demonstrated—using PCA and low-rank matrix completion (LRMC) tasks—the advantage of increasing BS for computational time via SFO complexity. Specifically, we observed that increasing BS yields a better minimizer (compared with a small constant BS) in term of achieving a small gradient norm and a shorter computational time (compared with a large constant BS). These findings are consistent with our theoretical analysis, which shows that increasing BS improves the SFO rate for reaching $\|G_{T}\|^{2}\leq\epsilon^{2}$ from $O(\epsilon^{-4})$ (under constant BS with critical batch size) to $O(\epsilon^{-2})$ (under increasing BS in the experimental setting).
@@ 2 Preliminaries
Let $\mathcal{M}$ be a Riemannian manifold. $T\mathcal{M}$ denotes the tangent bundle of $\mathcal{M}$. For all $x\in\mathcal{M}$, the inner product of $\mathcal{M}$ is represented $\langle\cdot,\cdot\rangle_{x}$ $T_{x}\mathcal{M}$ and induce the norm $\|\cdot\|_{x}$. For a smooth map $f\colon\mathcal{M}\rightarrow\mathbb{R}$, we can define gradient $\mathrm{grad}f$ as a unique vector field that satisfies $\forall v\in T_{x}\mathcal{M}\colon Df(x)[v]=\langle\mathrm{grad}f(x),v\rangle_{x}$. Although, in Euclidean space, iterative methods are updated by addition, Riemannian manifold does not equipped with addition generally. There is an alternative of updating procedure as follows.
Definition 1 (Retraction).
@@
Let $0_{x}$ be a zero element of $T_{x}\mathcal{M}$. When a map $R\colon T\mathcal{M}\ni(x,v)\mapsto R_{x}(v)\in\mathcal{M}$ satisfies the following conditions, $R$ is referred to as a retraction on $\mathcal{M}$.
  1. 1.
    $R_{x}(0_{x})=x$,
  2. 2.
    With the canonical identification $T_{0_{x}}T_{x}\mathcal{M}\simeq T_{x}\mathcal{M}$, $DR_{x}(0_{x})=\mathrm{id}_{T_{x}\mathcal{M}}$ for all $x\in\mathcal{M}$,
where $\mathrm{id}_{T_{x}\mathcal{M}}\colon T_{x}\mathcal{M}\to T_{x}\mathcal{M}$ is the identity mapping.
Iterative methods on Riemannian manifold are defined by $x_{t+1}=R_{x_{t}}(\eta_{t}d_{t})$ generally. Exponential maps are often used as retractions. The following assumption plays a central role in Lemma 1.
Assumption 1 (Retraction Smoothness).
Let $f\colon\mathcal{M}\rightarrow\mathbb{R}$ be a smooth map. Then there exists $L_{r}>0$ such that, for all $x\in\mathcal{M}$ and all $v\in T_{x}\mathcal{M}$,
$f(R_{x}(v))\leq f(x)+\langle\mathrm{grad}f(x),v\rangle_{x}+\frac{L_{r}}{2}\|v\|_{x}^{2}.$
In the Euclidean space setting, $L$-smoothness implies a property similar to retraction smoothness. The property corresponding to $L$-smoothness in Euclidean space is defined for $f\colon\mathcal{M}\rightarrow\mathbb{R}$ as
$\exists L>0,\forall x,y\in\mathcal{M}\colon\|\mathrm{grad}f(x)-\Gamma^{x}_{y}\mathrm{grad}f(y)\|_{x}\leq L_{r}d(x,y),$
@@
where $\Gamma$ is the parallel transport from $T_{y}\mathcal{M}$ to $T_{x}\mathcal{M}$, and $d(\cdot,\cdot)$ is the Riemannian distance. This is a sufficient condition for Assumption 1 with $R\coloneq\mathrm{Exp}$ (see Boumal (2023, Corollary 10.54)). This case is frequently used (e.g., Zhang & Sra (2016); Criscitiello & Boumal (2023); Kim & Yang (2022); Liu et al. (2017)). Other sufficient conditions for Assumption 1 were identified by Kasai et al. (2018, Lemma 3.5) or Sakai & Iiduka (2025, Proposition 3.2). Now, we consider the empirical risk minimization problem such that
$\underset{x\in\mathcal{M}}{\text{minimize}}\quad f(x)\coloneq\frac{1}{N}\sum_{j=1}^{N}f_{j}(x),$
@@
where every $f_{j}\colon\mathcal{M}\rightarrow\mathbb{R}$ is smooth and lower bounded. Therefore, $f$ satisfies the same conditions. This assumption is often made for both Euclidean space and Riemannian space. The lower boundedness of $f$ is essential for analyses using optimization theory since unbounded $f$ may not have optimizers. We denote a minimizer of $f$ as $x^{\star}$ ($f^{\star}\coloneq f(x^{\star})$) and denote $N$ is the size of a dataset. In many machine learning tasks, the dimension of model parameters $x$ or the size of dataset $N$ is large. Hence, we use the following mini-batch gradient to efficiently approximate the gradient of $f$.
$\mathrm{grad}f_{B}(x)\coloneq\frac{1}{b}\sum_{j=1}^{b}\mathrm{grad}f_{\xi_{j}}(x),$
where $B$ represents a minibatch with size $b$, and $(\xi_{j})_{j}$ is a $\{1,\cdots,N\}$-valued i.i.d. copy. Namely, the gradient of $f_{B}(x)=\frac{1}{b}\sum_{j=1}^{b}f_{\xi_{j}}(x)$ is used instead of $\mathrm{grad}f(x)$. The following assumption is reasonable given this situation.
Assumption 2 (Bounded Variance Estimator).
The stochastic gradient given by a distribution $\Pi$ is an unbiased estimator and has a bounded variance:
  1. 1.
    $\forall x\in\mathcal{M}\colon\mathbb{E}_{\xi\sim\Pi}\left[\mathrm{grad}f_{\xi}(x)\right]=\mathrm{grad}f(x)$,
  2. 2.
    $\exists\sigma>0,\forall x\in\mathcal{M}\colon\mathbb{V}_{\xi\sim\Pi}(\mathrm{grad}f_{\xi}(x))\leq\sigma^{2}.$
Remark 1.
Note that, from 1, $\mathbb{V}_{\xi\sim\Pi}(\mathrm{grad}f_{\xi}(x))=\mathbb{E}_{\xi\sim\Pi}\left[\|\mathrm{grad}f_{\xi}(x)-\mathrm{grad}f(x)\|_{x}^{2}\right]$ holds.
For example, if $f$ being lower bounded satisfies Assumption 1 and one takes the uniform distribution as $\Pi$, then Assumption 2 holds. RSGD is defined as
$x_{t+1}=R_{x_{t}}(-\eta_{t}\mathrm{grad}f_{B_{t}}(x_{t})).$
We suppose that the sequence $(x_{t})_{t}$ generated by RSGD is independent of random variables $(\xi_{j,t})_{j,t}$ for the minibatch gradient. Note that the BS $(b_{t})_{t}$ is variable at each step.

### v4

@@ Abstract
We theoretically analyzed the convergence behavior of Riemannian stochastic gradient descent (RSGD) and found that using an increasing batch size leads to faster convergence than using a constant batch size, not only with a constant learning rate but also with a decaying learning rate, such as cosine annealing decay and polynomial decay. The convergence rate improves from $O(T^{-1}+C)$ with a constant batch size to $O(T^{-1})$ with an increasing batch size, where $T$ denotes the total number of iterations and $C$ is a constant. Using principal component analysis and low-rank matrix completion, we investigated, both theoretically and numerically, how an increasing batch size affects computational time as quantified by stochastic first-order oracle (SFO) complexity. An increasing batch size was found to reduce the SFO complexity of RSGD. Furthermore, an increasing batch size was found to offer the advantages of both small and large constant batch sizes.
keywords:
Batch Size; Stochastic First-order Oracle Complexity; Learning Rate; Riemannian Optimization; Riemannian Stochastic Gradient Descent
@@ 1 Introduction
Study/Theorem Batch Size Learning Rate Convergence Analysis
Bonnabel (2013) s.a. LR $\underset{t\to\infty}{\lim}\mathbb{E}[\mathrm{grad}f(x_{t})]=0\text{ a.s.}$
Zhang and Sra (2016) $\eta_{t}=O((L_{r}+\sqrt{t})^{-1})$ $\mathbb{E}[f(x_{T})-f^{\star}]=O(\sqrt{T}(C+T)^{-1})$
Tripuraneni et al. (2018) Constant $\eta_{t}=O(t^{-\alpha})$ $\mathbb{E}\|\Delta_{T}\|=O(T^{-\frac{1}{2}})$
Hosseini and Sra (2020) Constant $\eta=O(T^{-\frac{1}{2}})$ $\|G_{T}\|^{2}=O(T^{-\frac{1}{2}})$
Sakai and Iiduka (2024) Constant $\eta_{t}=O((t+1)^{-\frac{1}{2}})$ $\|G_{T}\|^{2}=O(T^{-\frac{1}{2}}\log T)$
Sakai and Iiduka (2025) Increasing Constant $\|G_{T}\|^{2}=O(T^{-1})$
Theorem 3.2 Constant Constant and Decay $\|G_{T}\|^{2}=O(T^{-1}+C)$
Theorem 3.3 Increasing Constant and Decay $\|G_{T}\|^{2}=O(T^{-1})$
Theorem 3.4 Increasing Warm-up $\|G_{T}\|^{2}=O((T-T_{w})^{-1})$
Theorem 3.5 Constant Warm-up $\|G_{T}\|^{2}=O((T-T_{w})^{-1}+C)$
Table 1: Comparison of RSGD convergence analyses, where $\alpha\in(\frac{1}{2},1),C_{1}\in(0,1),C_{2}$ and $C$ are constants, and $(x_{t})$ is a sequence generated by RSGD. The usual stochastic approximation LR (s.a. LR) is defined as $\sum_{t=0}^{\infty}\eta_{t}=\infty,\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty$. $\mathbb{E}[\|\Delta_{T}\|]\coloneq\mathbb{E}[\|R_{x^{\star}}^{-1}(x_{T})\|]$. $T$ is the total number of iterations, and $T_{w}$ is the number of warm-up iterations (see Section 3.3). $\|G_{T}\|^{2}\coloneq\min_{t\in\{0,\cdots,T-1\}}\mathbb{E}[\|\mathrm{grad}f(x_{t})\|^{2}_{x_{t}}]$. ‘Decay’ includes LR schedules of the form $\eta_{t}=O(t^{-\frac{1}{2}})$. More detailed explanations of the convergence criteria are provided in Appendix F.
Stochastic gradient descent (SGD) (Robbins and Monro, 1951) is a basic algorithm, widely used in machine learning (Liu et al., 2021; He et al., 2016; Krizhevsky et al., 2012). Riemannian stochastic gradient descent (RSGD) was introduced in Bonnabel (2013). Riemannian optimization (Absil et al., 2008; Boumal, 2023), which addresses RSGD and its variants, has attracted attention due to its potential application in various tasks, including many machine learning tasks. For example, it has been used in principal components analysis (PCA) (Breloy et al., 2021; Liu and Boumal, 2020; Roy et al., 2018), low-rank matrix completion (Vandereycken, 2013; Boumal and Absil, 2015; Kasai and Mishra, 2016), convolutional neural networks (Wang et al., 2020; Huang et al., 2017), and graph neural networks (Zhu et al., 2020; Chami et al., 2019; Liu et al., 2019). It has also been used in applications of optimal transportation theory (Lin et al., 2020b; Weber and Sra, 2023). It is well established that the performance of Euclidean SGD strongly depends on both the batch size (BS) and learning rate (LR) settings (Goyal et al., 2017; Smith et al., 2018; Zhang et al., 2019; Lin et al., 2020a). In the context of Riemannian SGD, Ji et al. (2024); Bonnabel (2013); Kasai et al. (2019, 2018) addressed the use of a constant BS, and Sakai and Iiduka (2025); Han and Gao (2022) addressed the use of an adaptive BS, although the latter did not consider RSGD specifically. In parallel, various decaying LR strategies, including cosine annealing (Loshchilov and Hutter, 2017) and polynomial decay (Chen et al., 2018), have been devised, and their effectiveness has been demonstrated for the Euclidean case. Motivated by these previous studies, we developed a novel convergence analysis of RSGD, proved that the convergence rate is improved with an increasing BS, and showed that the stochastic first-order oracle (SFO) complexity is reduced with an increasing BS. Our numerical results support our theoretical results.
@@ 1.1 Previous Results and Our Contributions
Constraint: Bonnabel (2013); Zhang and Sra (2016); Durmus et al. (2021); Sakai and Iiduka (2024) considered Hadamard manifolds or their submanifolds. Tripuraneni et al. (2018) treated Riemannian manifolds that satisfy the condition required to ensure a unique geodesic connecting any two points, including Hadamard manifolds. Sakai and Iiduka (2025) addressed embedded submanifolds of Euclidean space. In contrast, we considered general Riemannian manifolds that encompass all of the above conditions.
@@
Learning rate: Bonnabel (2013) considered s.a. LRs, and Zhang and Sra (2016); Tripuraneni et al. (2018); Sakai and Iiduka (2024) considered diminishing LRs of the form $\eta_{t}=\frac{1}{t^{a}}$. Hosseini and Sra (2020); Durmus et al. (2021) considered constant LRs. Sakai and Iiduka (2025) used both constant and diminishing LRs. We considered constant, diminishing, cosine annealing, and polynomial decay LRs, as well as their warm-up versions.
@@
Batch size: Because Bonnabel (2013) considered expected risk minimization using expectation instead of sample mean—as is customary in empirical risk minimization—the concept of BS is not generally applicable in the expected-risk-minimization setting. Zhang and Sra (2016); Tripuraneni et al. (2018); Hosseini and Sra (2020); Durmus et al. (2021); Sakai and Iiduka (2024) considered a constant BS. Sakai and Iiduka (2025) and this study considered an increasing BS. Furthermore, we numerically and theoretically compared using a constant BS with using an increasing BS on the basis of SFO complexity. Our contributions are as follows: (i) we theoretically showed that using an increasing BS yields a more favorable SFO complexity of $O(\epsilon^{-2})$, whereas using a constant BS equal to the critical BS yields a rate of $O(\epsilon^{-4})$; (ii) we numerically observed that an increasing BS yields both a better optimal solution—in terms of attaining a smaller gradient norm—and a shorter computational time than either a small or a large constant BS (see Section 4.3 for more details).
@@
Objective function: Bonnabel (2013) considered three times continuously differentiable functions, with both the gradient and Hessian uniformly bounded. Tripuraneni et al. (2018) treated functions that are twice continuously differential, subject to additional conditions on the Hessian. Other studies considered only once continuously differentiable functions. Zhang and Sra (2016) addressed convex and smooth functions, while Durmus et al. (2021) considered strongly convex and smooth functions. Hosseini and Sra (2020); Sakai and Iiduka (2024, 2025) treated more general classes of functions, specifically nonconvex functions with bounded gradients. We treated nonconvex functions without bounded gradients.
@@
Convergence rate: A comparison of the previous and current RSGD convergence analyses is presented in Table 1; here we outline the key results. Tripuraneni et al. (2018) obtained a convergence rate of $O(T^{-\frac{1}{2}})$ (with a criterion other than the gradient norm) using an LR of $\eta=O(T^{-\frac{1}{2}})$; however, this rate decays to zero as $T\to\infty$, making it unsuitable for practical use. Durmus et al. (2021) attained exponential convergence under strong convexity and smoothness assumptions. Sakai and Iiduka (2025) attained a rate of $O(T^{-1})$ for several adaptive methods, including RSGD. Our work differs from Sakai and Iiduka (2025) in four key aspects : (i) we provide both theoretical and numerical evidence—based on SFO complexity—that using an increasing BS leads to better performance than using a constant BS; (ii) in addition to a constant LR and a diminishing LR, we addressed cosine annealing, polynomial decay, and warm-up LRs; (iii) Sakai and Iiduka (2024) focused on embedded submanifolds of Euclidean space unlike us; (iv) we do not assume bounded gradient norms, thereby expanding the range of applicable scenarios.
Contributions:
@@
  • We developed a convergence analysis of RSGD that incorporates an increasing BS, a cosine annealing LR, and a polynomial decay LR, under assumptions more general than those used in prior work. Our analysis shows that using an increasing BS improves the convergence rate of RSGD from $O(T^{-1}+C)$ to $O(T^{-1})$.
@@
  • We numerically and theoretically demonstrated—using PCA and low-rank matrix completion (LRMC) tasks—the advantage of an increasing BS in reducing computational time, as quantified by SFO complexity. Specifically, we observed that an increasing BS yields a better optimal solution (compared with a small constant BS) in terms of attaining a smaller gradient norm and a shorter computation time (compared with a large constant BS). These findings are consistent with our theoretical analysis, which shows that an increasing BS reduces the SFO complexity for achieving $\|G_{T}\|^{2}\leq\epsilon^{2}$ from $O(\epsilon^{-4})$—with a constant BS set equal to the critical BS—to $O(\epsilon^{-2})$ in the experimental setting.
@@ 2 Preliminaries
Let $\mathcal{M}$ be a Riemannian manifold and $T_{x}\mathcal{M}$ denote the tangent space at $x\in\mathcal{M}$. On $T_{x}\mathcal{M}$, the inner product of $\mathcal{M}$ is denoted by $\langle\cdot,\cdot\rangle_{x}$ and induces the norm $\|\cdot\|_{x}$. For a smooth map $f\colon\mathcal{M}\rightarrow\mathbb{R}$, we can define gradient $\mathrm{grad}f$ as a unique vector field that satisfies $\forall v\in T_{x}\mathcal{M}\colon Df(x)[v]=\langle\mathrm{grad}f(x),v\rangle_{x}$. A Riemannian manifold is generally not equipped with vector addition, whereas in Euclidean space iterative methods are updated by addition. Iterative updates are instead performed using alternative operations—such as retraction.
Definition 2.1 (Retraction).
@@
Let $0_{x}$ denote the zero element of $T_{x}\mathcal{M}$. A map $R\colon T\mathcal{M}\ni(x,v)\mapsto R_{x}(v)\in\mathcal{M}$ is called a retraction on $\mathcal{M}$ if it satisfies the two following conditions for all $x\in\mathcal{M}$. (I) $R_{x}(0_{x})=x$; (II) with the canonical identification $T_{0_{x}}T_{x}\mathcal{M}\simeq T_{x}\mathcal{M}$, $DR_{x}(0_{x})=\mathrm{id}_{T_{x}\mathcal{M}}$, where $\mathrm{id}_{T_{x}\mathcal{M}}\colon T_{x}\mathcal{M}\to T_{x}\mathcal{M}$ denotes the identity map.
@@
Iterative methods on Riemannian manifold are defined by $x_{t+1}=R_{x_{t}}(\eta_{t}d_{t})$ generally. Exponential maps are often used as retractions. The following assumption plays a central role in Lemma 3.1. {assumption} [Retraction Smoothness] Let $f\colon\mathcal{M}\rightarrow\mathbb{R}$ be a smooth map. Then there exists $L_{r}>0$, such that $^{\forall}x\in\mathcal{M},^{\forall}v\in T_{x}\mathcal{M}$, $f(R_{x}(v))\leq f(x)+\langle\mathrm{grad}f(x),v\rangle_{x}+\frac{L_{r}}{2}\|v\|_{x}^{2}$. In the Euclidean space setting, $L$-smoothness implies a property similar to retraction smoothness. The property corresponding to $L$-smoothness in Euclidean space is defined for $f\colon\mathcal{M}\rightarrow\mathbb{R}$ as $^{\exists}L>0,^{\forall}x,y\in\mathcal{M}\colon\|\mathrm{grad}f(x)-\Gamma^{x}_{y}\mathrm{grad}f(y)\|_{x}\leq L_{r}d(x,y)$, where $\Gamma$ is the parallel transport from $T_{y}\mathcal{M}$ to $T_{x}\mathcal{M}$, and $d(\cdot,\cdot)$ is the Riemannian distance. This condition is sufficient for Assumption 2.1 with $R\coloneq\mathrm{Exp}$ (see Boumal (2023, Corollary 10.54)). This case is frequently used (e.g., Zhang and Sra (2016); Criscitiello and Boumal (2023); Kim and Yang (2022); Liu et al. (2017)). Other sufficient conditions for Assumption 2.1 were identified by Kasai et al. (2018, Lemma 3.5) and Sakai and Iiduka (2025, Proposition 3.2). Now, we consider the empirical risk minimization problem such that
$\text{minimize}\ f(x)\coloneq\frac{1}{N}\sum_{j=1}^{N}f_{j}(x)\quad\text{subject to}\ x\in\mathcal{M},$
@@
where each $f_{j}\colon\mathcal{M}\rightarrow\mathbb{R}$ is smooth and lower bounded. Therefore, $f$ is also smooth and lower bounded. This assumption is often made for both Euclidean space and Riemannian space. The lower boundedness of $f$ is essential for analyses using optimization theory because unbounded $f$ may not have optimizers. We denote an optimal value of $f$ by $f^{\star}$ and let $N$ denote the size of the dataset. In many machine learning tasks, either the dimension of model parameters $x$ or the size of dataset $N$ is large. Hence, we use the following minibatch gradient to efficiently approximate the gradient of $f$: $\mathrm{grad}f_{B}(x)\coloneq\frac{1}{b}\sum_{j=1}^{b}\mathrm{grad}f_{\xi_{j}}(x)$, where $B$ represents a minibatch with size $b$, and $(\xi_{j})_{j=1}^{b}$ is a sequence of $\{1,\cdots,N\}$-valued i.i.d. random variables distributed as $\Pi$. That is, the gradient of $f_{B}(x)=\frac{1}{b}\sum_{j=1}^{b}f_{\xi_{j}}(x)$ is used instead of $\mathrm{grad}f(x)$. The following assumption is reasonable given this situation. {assumption} [Bounded Variance Estimator] The stochastic gradient given by a distribution $\Pi$ is an unbiased estimator of the full gradient and has bounded variance: (I) $\forall x\in\mathcal{M}\colon\mathbb{E}_{\xi\sim\Pi}\left[\mathrm{grad}f_{\xi}(x)\right]=\mathrm{grad}f(x)$; (II) $\exists\sigma>0,\forall x\in\mathcal{M}\colon\mathbb{V}_{\xi\sim\Pi}(\mathrm{grad}f_{\xi}(x))\leq\sigma^{2}$. For example, if $f$ is lower bounded, satisfies Assumption 2.1, and $\Pi$ is taken to be the uniform distribution over $\{1,\ldots,N\}$, then Assumption 2 holds. RSGD is defined by $x_{t+1}=R_{x_{t}}(-\eta_{t}\mathrm{grad}f_{B_{t}}(x_{t})).$ We assume that the sequence $(x_{t})_{t}$ generated by RSGD is independent of the random variables $(\xi_{j,t})_{j,t}$ used in the minibatch gradient. Note that the BS $(b_{t})_{t}$ may vary at each iteration.